OpenStudy (anonymous):

Determine the type of boundary line and shading for the graph of the inequality 5x + y greater than or equal to -10.

OpenStudy (anonymous):

Dashed line with shading on the side that includes the origin Solid line with shading on the side that does not include the origin Dashed line with shading on the side that does not include the origin Solid line with shading on the side that includes the origin

OpenStudy (texaschic101):

Solid line means there is an equal sign in the problem Dashed line means there is no equal sign in the problem The origin is (0,0)...if you sub this into your inequality and the inequality comes out true, then the shading includes the origin. If it is not true, the shading does not include the origin. lets begin... 5x + y >= -10 .... there is an equal sign, so it has a solid line (0,0) 5(0) + 0 >= -10 0 >= -10 (false) shading that does not include the origin conclusion : Solid line with shading on the side that does not include the origin

OpenStudy (anonymous):

if u graph it though, the shading is over (0,0)

OpenStudy (texaschic101):

wait...I messed up...lol..... (0,0) 5x + y >= -10 5(0) + 0 >= -10 0 >= -10 (true) solid line with shading that includes the origin The origin is (0,0)...x = 0 and y = 0 sub those points into the inequality and see if it makes the inequality true...if it does, it includes the origin

OpenStudy (anonymous):

so its d?

OpenStudy (texaschic101):

yep..its D

OpenStudy (anonymous):

thx, can u help me with another 1?

OpenStudy (texaschic101):

sure

OpenStudy (anonymous):

The following table shows the values of f(x) and g(x) for different values of x: Which of the following best describes the function f(x) + g(x)? a. 2x + 2 b. 2x + 2 c. 2x + 2x d. 2(2x)

OpenStudy (anonymous):

hold on i messed up on the multiple choise

OpenStudy (texaschic101):

I am sorry....I am not good with functions. You will need someone who knows more about this to help you. Again, I am sorry

OpenStudy (anonymous):

a. \[2^{x+2}\] b. \[2^{x}+2\] c. \[2^{x}+2x\] d. \[2(2^x)\]

OpenStudy (anonymous):

nvmd... im guessing

OpenStudy (anonymous):

I got it right!!!

OpenStudy (anonymous):

its b